[figure: 9.7 cm]

EFFECT OF LIGHT ON MAGNETIC STATE OF THE MANGANITE Pr_{0.6}La_{0.1}Ca_{0.3}MnO_{3} THIN FILMS

S.L. Gnatchenko^{ab}, M. Baran^{a}, O.Yu. Gorbenko^{c}, A.R. Kaul^{c}, R. Szymczak^{a}, H. Szymczak^{a} a Institute of Physics, Polish Academy of Sciences, Al. Lotników 32/46, 02-668 Warsaw, Poland

b Inst. for Low Temperature Phys. & Engineering, National Acad. of Sci. of Ukraine

47 Lenin Ave., 310164 Kharkov, Ukraine c Moscow State University, 119899 Moscow, Russia

Effect of illumination of He-Ne (λ = 633 nm) and Ar (λ = 488 nm) lasers on magnetization of manganite Pr_{0.6}La_{0.1}Ca_{0.3}MnO_{3} epitaxial thin films was studied by means of SQUID magnetometer equipped with Fiber Optic Sample Holder. It has been found that the magnetization of the films increases under illumination in a magnetic field near the field of phase transition from antiferromagnetic (AFM) to ferromagnetic (FM) state. The obtained experimental results allow to conclude that visible light stimulates this phase transition. The increase of magnetization is apparently related to a partial transformation of AFM phase into FM one and a formation of ferromagnetic clusters within the AFM state. A relaxation of the photoexcited state was not detected for one hour. Light-induced state of the manganite is stable or long-lived metastable one.

[figure: 13.4 cm]

Subject category :

2. Magnetic Films, Surfaces and Multilayers

Presentation mode : poster

Corresponding author : M. Baran

Address for correspondence : Dr M. Baran Instytut Fizyki PAN Al. Lotników 32/46 02-668 Warszawa

Email address : baran@ifpan.edu.pl